The diameter of an electric cable is normally distributed, with a mean of 0.8 inch and a standard deviation of 0.01 inch. What is the probability that the diameter will exceed 0.81 inch? (You may need to use the standard normal distribution table. Round your answer to three decimal places.)
Answer #1

P(X < A) = P(Z < (A - mean)/standard deviation)

Mean = 0.8 inch

Standard deviation = 0.01 inch

P(the diameter will exceed 0.81 inch) = P(X > 0.81)

= 1 - P(X < 0.81)

= 1 - P(Z < (0.81 - 0.8)/0.01)

= 1 - P(Z < 1)

= 1 - 0.8413

= 0.1587